• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Moffit

    Log entry time 12:24:27 on April10,2008

    Entry number 231179

    keyword=Access activity

    Re-routed the ethernet cable from the halladaq-cat switch (port 14) from the bblinux? switch to the DAQ weldment switch. Connected the bblinux? switch to port 13 of the halladaq-cat switch.

    Added some lead brick around both network switches.

    Plugged in a copy of the F1TDC reference to F1 TDC slot 6, ch33.

    Checked and refilled the Bigbite helium balloons.

    Found that the bigbite stand had shifted to a slightly larger angle (probably during a LHRS move, if they are slightly coupled at the pivot... or by movement/activity on the stand while working on the trigger planes/wirechambers). The slight change showed up as a 2 cm shift at the track.

    LHRS was moved to 12.5° by Scot.